# T-Explainer: A Model-Agnostic Explainability Framework Based on Gradients

### v2

@@ 3.2 T-Explainer: Computational Aspects
Figure 1 illustrates the step-by-step pipeline of T-Explainer for feature attribution. Computing the gradient of a known real-valued function is (theoretically) simple, demanding to take the partial derivatives of the function. However, we need to calculate partial derivatives of an arbitrary black-box model $f$, which was previously trained and holds complex internal mechanisms. To that end, we perturb the instance $\mathbf{x}$ attribute-wise and compute the respective change in the model’s output, approximating the partial derivatives through centered finite differences [36]:
@@
Finite differences are well-established approaches to approximate differential equations, replacing the derivatives with discrete approximations (such transformation results in computationally feasible systems of equations) [36].
Specifically, we rely on the centered finite difference because that approach simply averages two one-sided perturbations of each attribute, resulting in a second-order accurate approximation with an error proportional to $||\mathbf{h}||^{2}$ [36]. The centered finite difference approximation of the partial derivatives demands to set up the magnitude of the perturbation $\mathbf{h}$. The displacement $\mathbf{h}$ must be small enough to generate a perturbation close to the input data $\mathbf{x}$. By definition, the derivatives of $f$ are computed by making $\mathbf{h}\to 0$ in Equation 9. In other words, we have to set $\mathbf{h}$ to a small value.
In practice, if $\mathbf{h}$ is too small, it can generate significant round-off errors or singularity cases. On the other hand, if $\mathbf{h}$ is too large, it can lead to truncation errors and inappropriate approximations. As far as we know, there is no closed solution to determine an optimum value to $\mathbf{h}$. To solve this issue, we developed a $\mathbf{h}$ optimizer method based on a binary search that minimizes a mean squared error (MSE) cost function. The method runs over the $[h_{min},h_{max}]$ interval, searching for an optimal $\mathbf{h}$ to produce a good approximation of $f(\mathbf{x})$ from Equation 2. We set $h_{min}$ with the minimum distance between any two instances on the dataset, i.e., $h_{min}=\min_{i,j=1}^{n}||{\mathbf{x}_{i}}-{\mathbf{x}_{j}}||$, $\forall{i}\neq{j},\mathbf{x}_{i},\,\mathbf{x}_{j}\in\mathbf{X}$. The optimization method is formulated as follows:
@@
where $\mathcal{L}$ is the cost function (MSE in our case), $\nabla f_{\mathbf{x}}$ represents the approximated value of the gradient $\nabla f(\mathbf{x})$ in each iteration of the $\mathbf{h}$ optimizer algorithm computing the method in Equation 9 on Equation 2, $\epsilon_{\mathbf{x}}$ is the cost threshold upper-bounded by $(h_{min})^{2}$ [36], and $\theta$ is a method that ensures the numerical stability of the optimization process by checking if $\mathbf{h}$ is keeping the gradient (Jacobian matrix) $\nabla f_{\mathbf{x}}$ non-singular (or full-rank). It is important to highlight that local perturbations must preserve the data’s normalization range to avoid model extrapolations [30]. Further details about the optimization procedure are given in the Appendix Section. This approach can find the gradient ${\nabla f(\mathbf{x})}$ of a binary-class model or even the entire Jacobian related to all the model’s output probabilities. Once the gradient is estimated in $\mathbf{x}$, the importance of each attribute is computed as $\phi_{i}=\frac{\partial{f}(\mathbf{x})}{\partial x_{i}}$.
The explanations must remain consistent with the model’s behavior in the instance’s locality to ensure meaningful interpretability. Through the $\mathbf{h}$ optimization process, T-Explainer attributes feature importance using Taylor’s expansion for each instance tailored to its local characteristics. In other words, the T-Explainer is primarily designed as a local explainer. According to Ribeiro et al. [47], simultaneously achieving local and global fidelity in XAI is challenging because aggregating local explanations to estimate global feature importance might be ineffective due to local explanations being instance-specific and often inconsistent with global explanations. However, the adaptive $\mathbf{h}$ optimization strategy of T-Explainer is flexible enough to be extended and generate aggregated views while preserving local relationships, enhancing consistent and interpretable explanations locally and also at global levels.
@@ 3.2.1 Handling Categorical Data
All the considerations outlined above assume the attributes are continuous numerical data. However, most datasets are not limited to numerical data but also encompass categorical (nominal or qualitative) features. Handling categorical data is challenging, as many learning models cannot directly process nominal features, demanding numerical conversions. Different approaches can be applied to encode categorical values as numerical representations, including the well-known one-hot encoding [49].
Numerical encodings introduce challenges for XAI methods, particularly for the gradient-based ones. Nominal values represented with one-hot encode become binary constant values (zeros and ones), thus being discontinuous attributes where partial derivatives can not be properly estimated, inducing the explainer to erroneously attribute null importance to features that may significantly impact the prediction.
T-Explainer has a modular design that enhances flexible additions or improvements of functionalities. To address the challenges of categorical features, we implemented a mechanism to handle one-hot encoded columns into the T-Explainer framework. Once the user identifies the one-hot encoded categorical attributes, a procedure transforms them into intervals through continuous perturbations. Specifically, the 00 and $1$ values (resulting from one-hot encoding) are uniformly perturbed with displacement in the interval $\delta\in(-0.5,0.5)$, creating a range of values around 00 and $1$, simulating a continuous set of values around these two values.
The continuity induction procedure is performed over a copy of the training dataset, with the perturbed columns being normalized to the same interval of the numerical features (typically in the $[0,1]$ range). After that, transfer learning is applied in a copy of the original model, fitting it with the perturbed training dataset (the categorical fitted model). Finally, the T-Explainer runs on the newly trained model. In summary, predictions of instances holding categorical features are explained using the T-Explainer’s core implementation. However, we incorporated a preliminary layer in T-Explainer’s pipeline to handle categorical data.
The perturbation radius $\delta$ can impact the accuracy of the categorical fitted model, making it disagree with the original model. However, the disagreement varies depending on the training data. In our experiments, the proposed transformation of categorical attributes to numerical ones does not significantly impact the accuracy of the categorical fitted model for perturbations into small intervals. We fixed $\delta=0.1$ as the perturbation radius, which ensured the accuracy preservation.
@@ 4 Quantitative Metrics
A common drawback of XAI is the lack of quantitative evaluation of XAI methods. Most proposals rely on visual inspections or simplified human-centered case studies to check whether an explanation “looks reasonable.” However, such anecdotal strategies are prone to subjectivities and do not allow for formal comparisons between explanation methods, which is insufficient to ensure a robust verification of explanations’ consistency, continuity, and correctness [42]. This Section describes the evaluation metrics integrated into the T-Explainer framework. We focused our evaluations on automated quantitative metrics that enable us to compare different explanation methods regarding their stability, faithfulness, and local accuracy preservation.
A stability metric measures the robustness or sensitivity of an explainer when exposed to slightly different versions of an original input sample. Relative Input Stability (RIS) and Relative Output Stability (ROS) [3, 4] are metrics used to evaluate the stability of local explanations to changes in input data and output prediction probabilities, respectively. Alvarez-Melis and Jaakkola [6] introduced the stability concept, which Agarwal et al. [3] recently enhanced through the “relative stability” formulation. We followed theoretical definitions of RIS and ROS as defined by Agarwal et al. [4]; however, we introduced improvements to address some identified issues, which are detailed below.
@@
$\forall\mathbf{x}^{\prime}\in\mathcal{N}_{\mathbf{x}}$, with ${p}$ defining the $l_{p}$ norm used to measure the changes and $\epsilon_{c}>0$ as a small values threshold to avoid zero division. The larger the RIS or ROS values, the more unstable the method is related to input or output prediction perturbation.
RIS and ROS handle values close to zero in normalization processes. However, the original versions implemented in Agarwal et al. [4] apply a clipping method that ignores negative values by trimming them to the small values threshold. Consequently, significant negative importance values are converted to a fixed threshold, leading the metrics to ignore instability in negative attributions. Then, we designed a clipping method that preserves significant values, either positive or negative, while discarding those close to zero.
Furthermore, the original implementations of RIS and ROS may also ignore instances near the model’s decision boundaries. We solved that issue by sampling perturbations of each $\mathbf{x}$, sorting the perturbed data based on the distance to $\mathbf{x}$, and generating a neighborhood holding a minimum set of perturbed instances. This way, we ensure coverage and avoid measurement omissions or potential crashes in the metrics. Additionally, we extended RIS and ROS to evaluate mean changes in explanations.
To complement our stability evaluation, we introduce the Run Explanation Stability (RES) metric to assess the stability of local explainers for multiple runs on non-perturbed inputs. The principle behind RES is simple – it generates multiple explanations for the same original input under fixed model and explanation method settings and measures the extent to which such explanations change. According to the consistency property, identical inputs have identical explanations [42]. Given $n$ explanations $e_{\mathbf{x}i},\dots,e_{\mathbf{x}n}$ of a same non-perturbed instance $\mathbf{x}$, let $\bar{e}_{\mathbf{x}}$ be the mean explanation from those $n$ explanations, RES is defined as:
@@
Although we are introducing RES as part of the T-Explainer framework contributions, the rationale behind it was inspired by the Reiteration Similarity metric [7]; however, we simplified the formulation by using standard deviation as the similarity measure.
RIS, ROS, and RES metrics are unitless quantifiers with no “ideal” desirable values. Reasonable values depend on the context of the application, meaning those metrics must be interpreted relatively by comparing their results across XAI methods [5]. However, lower values indicate methods with higher stability rates. Furthermore, RIS and ROS fit in the stability for slight variations strategy of the continuity category of XAI assessment, while RES fits in the consistency category [42].
A faithfulness metric evaluates the extent to which an explanation is faithful to the predictive model it explains. In this sense, we implemented the Prediction Gap on Important Features (PGI) metric to assess the methods’ faithfulness under the single/incremental deletion strategy for correctness [42]. Petsiuk et al. [44] introduced the intuition behind deletions, which Dai et al. [22] recently enhanced in the PGI metric. Given $top(k,e_{\mathbf{x}})$ being the $k$ most important features determined by a local explanation $e_{\mathbf{x}}$, we implemented PGI iterating on $m$ input instances as follows:
$\text{PGI}(\mathbf{x},f,e_{\mathbf{x}})=\frac{1}{m}\sum_{j=1}^{m}[1-|f(\mathbf{x})-f(\mathbf{\tilde{x}}_{j})|]$ (14)
where ${\tilde{x}_{i}}=0[i\notin top(k,e_{\mathbf{x}})]$ and $f$ returns predicted probabilities. PGI values will be in the $[0,1]$ range; the higher the value, the more faithful the explanations are compared to the “true” black-box behavior.
We also introduce the Local Accuracy Preservation (LAP) metric, which assesses the extent to which feature importance explainers preserve the local accuracy property, i.e., the model prediction should be reconstructed by the summation of the importance values (see Equation 5). Given an output prediction probability $f(\mathbf{x})$ and its respective explanation $g_{\mathbf{x}}$, we define the LAP metric as the frequency ratio in which $|f(\mathbf{x})-g_{\mathbf{x}}|\leq\epsilon$, $\forall\mathbf{x}\in\mathbf{X}$, with $\epsilon$ representing a tolerance error and $g$ following the additive modeling defined in Equation 4 [39]. The tolerance error is necessary since we do not expect any explanation modeling $g_{\mathbf{x}}$ to achieve a perfect match to the original model $f(\mathbf{x})$. LAP returns values in the $[0,1]$ range; the higher the value, the more faithful the explainer is regarding local accuracy preservation, i.e., the more an explainer keeps $g_{\mathbf{x}}$ close to $f(\mathbf{x})$, $\forall\mathbf{x}\in\mathbf{X}$, the more “locally accurate” it is.
Unlike previous implementations, we developed our metrics to run all the explainers under evaluation into the same execution cycles. It is an advantage regarding fair comparisons since it ensures the explainers are exposed to the same conditions. Our evaluations are based on five automated quantitative metrics, enabling users to compare T-Explainer’s performance with state-of-the-art feature importance methods formally. The metrics selected evaluate the consistency, continuity, and correctness dimensions of XAI evaluation [42], providing robust experimental results that go beyond the typical evaluations in XAI literature.
@@ 5.1 Experimental Setup
We trained Neural Networks from scikit-learn (https://scikit-learn.org/) and Gradient-boosted Tree Ensemble classifiers (Random Forests-based) from the XGBoost gradient boosting library [20] as the black-box models used throughout the experiments. Logistic Regressions and SVMs could also be viable alternatives, but we focused on Neural Networks and Random Forests due to their wide adoption in Machine Learning. The models were trained with a dataset split of $80\%$ for training and $20\%$ for testing. Such division was achieved using the $train\_test\_split$ method from scikit-learn, with a consistent shuffling seed across all datasets. We conducted a Grid Search to determine the hyperparameters that optimize the models’ performance.
@@
We selected three different Neural Network models from the hyperparameter tuning: a three- and a five-hidden-layer neural network holding $64$ neuron units per layer (3H-NN and 5H-64-NN, respectively) and a five-layer neural network with $[64,128,128,128,64]$ neurons in each layer (5H-128-NN). All the neural network models use a learning rate of $0.01$, alpha $0.0001$, and maximum training epochs of $500$. The selected neural networks achieved better performances considering our classification tasks. We based our architectures on previous works [10, 16], with extra neural units and additional layers significantly increasing the training time without noticeably increasing performance.
Following our hyperparameters tunning, the Tree Ensemble classifiers (XRFC) have $500$ estimators, a maximum depth of $6$, a learning rate of $0.01$, and a gamma equal to $1$. All the other hyperparameters are the libraries’ default for Neural Networks and Tree Ensembles. Those models are used as the base black boxes for comparing T-Explainer against SHAP [39], LIME [47], and three gradient-based methods from the Captum library [31] – Integrated Gradients, Input $\times$ Gradient, and DeepLIFT. Specifically, we used the SHAP explainer for neural networks and the TreeSHAP explainer [40] for tree-based classifiers. Although SHAP, LIME, and Gradient-based methods were proposed a few years back, we benchmark T-Explainer with them because they continue to be the most widely used feature attribution explainers in research and practice [4].
@@
To avoid round-off issues and enhance readability, we use scientific notation to represent values larger than ${10}^{5}$, while values smaller than ${10}^{-10}$ will be taken as zero.

### v3

@@ 3.2 T-Explainer: Computational Aspects
Figure 1 illustrates the step-by-step pipeline of T-Explainer for feature attribution. Computing the gradient of a known real-valued function is (theoretically) simple, demanding that the partial derivatives of the function be taken. However, we need to calculate the partial derivatives of an arbitrary black-box model $f$, which was previously trained and has complex internal mechanisms. To that end, we perturb the instance $\mathbf{x}$ attribute-wise and compute the respective change in the model’s output, approximating the partial derivatives through centered finite differences [27]:
@@
Finite differences are well-established approaches to approximate differential equations, replacing the derivatives with discrete approximations (such transformation results in computationally feasible systems of equations) [27]. Specifically, we rely on the centered finite difference because that approach averages two one-sided perturbations of each attribute, resulting in a second-order accurate approximation with an error proportional to $||\mathbf{h}||^{2}$ [27]. The centered finite difference approximation of the partial derivatives demands establishing the magnitude of the perturbation $\mathbf{h}$. The displacement $\mathbf{h}$ must be small enough to generate a perturbation close to the input data $\mathbf{x}$. By definition, the derivatives of $f$ are computed by making $\mathbf{h}\to 0$ in Equation 9. In other words, we have to set $\mathbf{h}$ to a small value.
In practice, if $\mathbf{h}$ is too small, it can generate significant round-off errors or singularity cases. However, if $\mathbf{h}$ is too large, it can lead to truncation errors and inappropriate approximations. To our knowledge, there is no closed solution to determine an optimum value of $\mathbf{h}$. To solve this issue, we developed a $\mathbf{h}$ optimizer method based on a binary search that minimizes a mean squared error (MSE) cost function. The method runs over the $[h_{min},h_{max}]$ interval, searching for an optimal $\mathbf{h}$ to produce a good approximation of $f(\mathbf{x})$ from Equation 2. We set $h_{min}$ with the minimum distance between any two instances on the dataset, i.e., $h_{min}=\min_{i,j=1}^{n}||{\mathbf{x}_{i}}-{\mathbf{x}_{j}}||$, $\forall{i}\neq{j},\mathbf{x}_{i},\,\mathbf{x}_{j}\in\mathbf{X}$. The optimization method is formulated as follows.
@@
where $\mathcal{L}$ is the cost function (MSE in our case), $\nabla f_{\mathbf{x}}$ represents the approximated value of the gradient $\nabla f(\mathbf{x})$ in each iteration of the $\mathbf{h}$ optimizer algorithm computing the method in Equation 9 , $\epsilon_{\mathbf{x}}$ is the upper limit of the cost threshold bound by $(h_{min})^{2}$ [27], and $\theta$ is a method that ensures the numerical stability of the optimization process by checking if $\nabla f(\mathbf{x})$ comes from a non-singular (or full-rank) Jacobian matrix.
In the presence of a singular matrix, perturbation methods such as finite differences suffer from numerical instability. This instability can lead to inaccurate approximations of the Jacobian, making it unreliable for further calculations or updates. In addition, a singular or rank-deficient Jacobian indicates that some input features or dimensions are linearly dependent, and this is problematic because it means that the model is not fully sensitive to certain input variations, leading to poor sensitivity analysis. Specific to the context of our optimizer, a singular Jacobian leads to a failure in convergence, propagating errors that lead the explainer to output one-hot-like feature importance vectors. Therefore, $\theta$ rejects $\mathbf{h}$ when it results in a singular (or rank-deficient) Jacobian.
It is important to highlight that local perturbations must preserve the normalization range of the data to avoid model extrapolations [21]. Further details about the optimization procedure are given in the Appendix. This approach can find the gradient ${\nabla f(\mathbf{x})}$ of a binary-class model or even the entire Jacobian related to all the model’s output probabilities. Once the gradient is estimated at $\mathbf{x}$, the importance of each attribute is computed as $\phi_{i}=\frac{\partial{f}(\mathbf{x})}{\partial x_{i}}$.
The explanations must remain consistent with the model’s behavior in the instance’s locality to ensure meaningful interpretability. Through the $\mathbf{h}$ optimization process, T-Explainer attributes feature importance using Taylor’s expansion for each instance tailored to its local characteristics. In other words, the T-Explainer is designed primarily as a local explainer. According to Ribeiro et al. [36], simultaneously achieving local and global fidelity in XAI is challenging because the aggregation of local explanations to estimate global feature importance could be ineffective. After all, local explanations are instance-specific and often inconsistent with global explanations. However, the adaptive $\mathbf{h}$ optimization strategy of T-Explainer is flexible enough to be extended and generate aggregated views while preserving local relationships, enhancing consistent and interpretable explanations at local and global levels.
@@ 3.2.1 Handling Categorical Data
All the considerations outlined in the T-Explainer’s theoretical foundations (see The T-Explainer Section) assume that the attributes are continuous numerical data. However, most datasets are not limited to numerical data but also encompass categorical (nominal or qualitative) features. Handling categorical data is challenging, as many learning models cannot directly process nominal features, demanding numerical conversions. Different approaches can be applied to encode categorical values as numerical representations, including the well-known one-hot encoding [37].
Numerical encodings introduce challenges for XAI methods, particularly those based on gradients. Nominal values represented with one-hot encode become binary constant values (zeros and ones), thus being discontinuous attributes where partial derivatives can not be properly estimated, inducing the explainer to erroneously attribute null importance to features that may significantly impact the prediction.
T-Explainer has a modular design that enhances flexible additions or improvements in functionalities. To address the challenges of categorical features, we implemented a mechanism to handle one-hot encoded columns into the T-Explainer framework. Once the user identifies the one-hot encoded categorical attributes, a procedure transforms them into intervals through continuous perturbations. Specifically, the values 00 and $1$ (resulting from one-hot encoding) are uniformly perturbed with displacement in the interval $\delta\in(-0.5,0.5)$, creating a range of values around 00 and $1$, simulating a continuous set of values around these two values.
The continuity induction procedure is performed on a copy of the training dataset, with the perturbed columns normalized to the same interval of the numerical features (typically in the $[0,1]$ range). After that, transfer learning is applied to a copy of the original model, fitting it with the perturbed training dataset (the categorical fitted model). Finally, the T-Explainer runs on the newly trained model. In summary, predictions of instances holding categorical features are explained using the T-Explainer’s core implementation. However, we incorporated a preliminary layer in T-Explainer’s pipeline to handle categorical data.
The perturbation radius $\delta$ can affect the accuracy of the categorical fitted model, making it disagree with the original model. However, the disagreement varies according to the training data. In our experiments, the proposed transformation of categorical attributes to numerical ones does not significantly impact the accuracy of the categorical fitted model for perturbations in small intervals. We fixed $\delta=0.1$ as the perturbation radius, which ensured accuracy preservation.
@@ 4 Quantitative Metrics
A common drawback of research proposing XAI methods is a lack of clarity on the use of quantitative evaluations [33]. Most proposals rely on visual inspections or simplified human-centered case studies to check whether an explanation “looks reasonable.” However, such anecdotal strategies are prone to subjectivities and do not allow for formal comparisons between explanation methods, which is insufficient to ensure a robust verification of explanations’ consistency, continuity, and correctness [33]. Evaluation in XAI evolved significantly in the past few years, and this Section describes the evaluation metrics integrated into the T-Explainer framework.
A stability metric measures the robustness or sensitivity of an explainer when exposed to slightly different versions of an original input sample. Relative Input Stability (RIS) and Relative Output Stability (ROS) [3] are metrics used to evaluate the stability of local explanations to changes in input data and output prediction probabilities, respectively, based on the “relative stability” formulation that applies normalized distances allowing direct comparisons between different feature importance methods. We followed the theoretical definitions of RIS and ROS defined by Agarwal et al. [3].
@@
$\forall\mathbf{x}^{\prime}\in\mathcal{N}_{\mathbf{x}}$, $\hat{y}_{\mathbf{x}}=\hat{y}_{\mathbf{x}^{\prime}}$, with ${p}$ defining the $l_{p}$ norm used to measure the changes and $\epsilon_{c}>0$ as a small value threshold to avoid zero division. The higher the RIS or ROS values, the more unstable the method is related to input or output prediction perturbation.
We identified some issues that prevented us from using the implementations of Agarwal et al. [3]. Notably, their framework lacks flexibility for incorporating novel explainers beyond those included in OpenXAI, and the separate execution of RIS and ROS metrics reduces their comparability due to differences in the perturbed samples drawn from normal distributions. In contrast, our T-Explainer framework offers a flexible Python module that can be readily extended to assess multiple feature attribution methods while computing RIS and ROS in parallel.
Previously, the selection of perturbed samples satisfying $\hat{y}{\mathbf{x}}=\hat{y}{\mathbf{x}^{\prime}}$ was performed randomly, with no guarantee that a sufficient number of samples would be retained – particularly near the model’s decision boundaries. We solved this limitation by sampling perturbations for each $\mathbf{x}$ and retaining that $\hat{y}_{\mathbf{x}}=\hat{y}_{\mathbf{x}^{\prime}}$ according to their distance to $\mathbf{x}$ [36], ensuring a neighborhood holding a minimum set of samples.
RIS and ROS handle values close to zero in normalization processes. Both large negative and positive values indicate high importance – the sign indicates the direction of the feature’s contribution to the prediction. Therefore, we designed a double-sided clipping method that preserves significant values, either positive or negative, while discarding only those close to zero. In addition, we extended RIS and ROS to provide mean stability values.
To complement our stability evaluation, we introduce the Run Explanation Stability (RES) metric to quantify the stability of local explainers across multiple runs on non-perturbed inputs. The core idea behind RES is straightforward: Under fixed model and explanation method settings, the same original input should yield identical explanations [33]. In practice, we generate $n$ explanations, $e_{\mathbf{x}1},\dots,e_{\mathbf{x}n}$, for the same original instance $\mathbf{x}$ and compute their standard deviation, returning the maximum deviation in a set of instances. The RES metric is defined as follows:
@@
Although we introduce RES as part of the T-Explainer framework, the theoretical foundation behind it came from the Reiteration Similarity metric [5]; however, we simplified the formulation by using the standard deviation as the similarity measure.
RIS, ROS, and RES metrics are unitless quantifiers with no “ideal” desirable values. Reasonable values depend on the context of the application, meaning those metrics must be interpreted relatively by comparing their results across XAI methods [4]. However, lower values indicate methods with higher stability rates. Therefore, RIS and ROS fit into the stability for slight variations strategy of the continuity category of XAI assessment, while RES fits in the consistency category [33].
A faithfulness metric evaluates the extent to which an explanation is faithful to the predictive model it explains. In this sense, we implemented the Prediction Gap on Important Features (PGI) [35, 14] metric to evaluate the methods’ faithfulness under the single/incremental deletion strategy for correctness [33]. Given $top(k,e_{\mathbf{x}})$ being the $k$ most important features determined by a local explanation $e_{\mathbf{x}}$, we implemented PGI iterating on $m$ input instances as follows:
$\text{PGI}(\mathbf{x},f,e_{\mathbf{x}})=\frac{1}{m}\sum_{j=1}^{m}[|f(\mathbf{x})-f(\mathbf{\tilde{x}}_{j})|]$ (14)
where ${\tilde{x}_{i}}=0\,\,\,[i\in top(k,e_{\mathbf{x}})]$ and $f$ return predicted probabilities. PGI values will be in the $[0,1]$ interval; the higher the value, the more faithful the explanations are compared to the “true” black-box behavior, as perturbing those $k$ features caused a significant change in prediction.
Unlike previous implementations, we developed metrics to run all the explainers under evaluation in the same execution cycles. It is an advantage regarding fair comparisons since it ensures the explainers are exposed to the same conditions. Our evaluations are based on automated quantitative metrics, enabling users to formally compare the performances of state-of-the-art feature importance methods. The selected metrics evaluate the consistency, continuity, and correctness dimensions of XAI evaluation [33], providing robust experimental results beyond the typical evaluations in the XAI literature.
@@ 5.1 Experimental Setup
We trained Neural Networks from scikit-learn library [^2] https://scikit-learn.org/ and Gradient-boosted Tree Ensembles from the XGBoost library [^3] https://xgboost.readthedocs.io/ as the black-box models used throughout the experiments. Logistic Regressions and SVMs could also be viable alternatives, but we focused on Neural Networks and Random Forests due to their wide adoption in Machine Learning. The models were trained with a dataset split of $80\%$ for training and $20\%$ for testing. Such division was achieved using the $train\_test\_split$ method from scikit-learn, with a consistent shuffling seed across all datasets. We conducted a Grid Search to determine the hyperparameters that optimize the models’ performance.
@@
We selected three different neural networks from hyperparameter tuning: Three- and five-layer neural networks that hold $64$ neuron units per layer (3H-NN and 5H-64-NN, with 9.7K and 18K trainable parameters, respectively) and a five-layer neural network with $[64,128,128,128,64]$ neurons in each layer (5H-128-NN) and roughly 51K trainable parameters. All neural network models use a learning rate of $0.01$, alpha $0.0001$, and maximum training epochs of $500$. The selected neural networks achieved better performances considering our classification tasks. We based our architectures on previous works [8, 12], with extra neural units and additional layers significantly increasing the training time without noticeably increasing performance.
The Tree Ensemble classifiers (XRFC) have $500$ estimators, a maximum depth of $6$, a learning rate of $0.01$, and a gamma equal to $1$. All the other hyperparameters are the libraries’ default for Neural Networks and Tree Ensembles. Those models are used as the base black boxes for comparing the T-Explainer against SHAP [30], LIME [36], and three gradient-based methods from the Captum library [22] – Integrated Gradients, Input $\times$ Gradient, and DeepLIFT. Specifically, we used the SHAP explainer for neural networks and the TreeSHAP explainer [31] for tree-based classifiers. Although SHAP, LIME, and Gradient-based methods were proposed a few years back, we benchmark T-Explainer with them because they continue to be the most widely used feature attribution explainers in research and practice [3].
@@
To improve readability and avoid roundoff issues, we use scientific notation to represent values larger than ${10}^{5}$, while values smaller than ${10}^{-10}$ will be taken as zero.